User name LeRose
Log entry time 15:54:58 on October31,2001
Entry number 75501
Arun Saha reports that the OTR should be considered non-functional. There's a problem with the filters. New parts are on order. It should be fixed next Tuesday. Contact Jean-Claude Denard for more information.